Do the functions 

y1(t)=√t and y2(t)=1/t

form a fundamental set of solutions of the 

equation 2t²y"+3ty'-y=0

, on the interval 

0 < t <infinity

Justify your answer.


1
Expert's answer
2021-05-31T16:19:31-0400

"y_1(t)={\sqrt t}"


"y_2(t)=\dfrac{1}{t}"


Given equation is -


"2t^{2}y''" "+" "3ty'-y=0"


Now to check "y_1(t)" and "y_2(t)" are set of solutions of this equation , we will check this by the help of wronskian Abel's theorem of method of solving differential equation .


"y_1(t)={\sqrt t}"


"y_1'(t)=\dfrac{1}{2{\sqrt t}}"

"y_1''(t)="


"y_2(t)=\dfrac{1}{t}"


"y_2'(t)=-\dfrac{1}{t^{2}}"

"y_2''(t)=2t^{-3}"


Now for checking that these two are the solution of this differential equation , we compute wronskian method of solving differential equation ,which is -


"\begin{vmatrix}
 y_1& y_2 \\
 y_1' & y_2'
\end{vmatrix}" ="y_1y_2'-y_2y_1'=\dfrac{3}{2}t^{\dfrac{-3}{2}}"


Notice than wronskian is zero at t=0 but non zero at t=1 . By the above corollary , "y_1" and "y_2 \" cannot be both be solution in the respective interval "0<t<{\infty}" .


Now to see whether given functions satisfy this diiferential equation we can also check it by expressing in terms of determinant -


"\begin{vmatrix}
 y_1 & y_2 &y\\
 y_1'& y_2' &y' \\
 y1'' &y_2'' & y''
 
\end{vmatrix}" "=0"



Now putting the values in determinant , and expanding along the 1st column , we get -


"\begin{vmatrix}
 {\sqrt t} & \dfrac{1}{t} &y\\
 \dfrac{1}{2{\sqrt t}}& \dfrac{-1}{t^{2}} &y' \\

 \dfrac{-1}{4}t^{\dfrac{-3}{2}} &2t^{-3} & y''
 
\end{vmatrix}" "=0"



Now expanding this determinant , we get -


"=2y''t^{2}-5ty'-5y=0"


which is form differential equation , so the given roots are not equation of the given equation .hence proved .